find speed if distance =10m and time =2s

Grade:

3 Answers

mohan kumar
11 years ago

We know that speed is distance/time

s=d/t

s=d/t

=10/2

=5

Therefore speed is 5m/s......

Thats all !!!!!!!!

Madhukar Thalore
11 years ago

Wek now speed =distance /time

So here speed=10/2=5m/s.

Its very easy.

veeravalli mohan
11 years ago

the formula for speed is distance travelled by time

speed = distance/time

here distance is 10m, time is 2s

so speed is 10/2 =5m/s